OpenStudy (anonymous):

Word problem: a cruise ship has 680 rooms with a revenue $92,500.00. Rooms with a view cost $160 per night and rooms without a view costs $105per night. how many rooms are there of each

OpenStudy (anonymous):

OK let's get you started... Let v = # of cabins with a view w = # of cabins with 4 walls. Can you create an equation using the 680 data point?

OpenStudy (anonymous):

I just don't understand

OpenStudy (anonymous):

hi, i will solve all your problems if you give me lifesaver

OpenStudy (anonymous):

i need back scratched,

OpenStudy (anonymous):

what is a life saver, this is my first time on this site

OpenStudy (anonymous):

click on my name become a fan

OpenStudy (anonymous):

OK we know that the sum of view and nonview cabins equals 680. <<-- This is equation 1. The next equation is the tough one. Let's break-it down: How much money do all the View cabins earn? How much do the wall cabins earn? For equation 2 you'll write and expression that adds those together to get a total revenue of $92,500,000.00

OpenStudy (anonymous):

Here's a hint: Eqn 1: __ v + __ w = 680 Eqn 2: (???)*v +(???)*w = 92500000.00

OpenStudy (anonymous):

X+y=680 160x+105y=92500.00 =y=680-x 160x+105(680-x)=160x+71400-105X=92500.00 =55x=21100 x=384 I am doing this correctly?

OpenStudy (anonymous):

Yes - good job. Now use "384" and substitute into equation #1 to find the number of wall-only cabins. Make sure you label your answer.

OpenStudy (anonymous):

384+y=680 680-384=296 y=296 Yes?

OpenStudy (anonymous):

Looks good - obviously it works in the first equation. To be extra safe you can plug both values back into the second to check your answer. Label you numbers! <<--Makes the teacher happy!

OpenStudy (anonymous):

THank you Thank You!!!! your are the bomb!!!

OpenStudy (anonymous):

Good luck!